Directrix (directrix):

Your opinion requested: Is this a well written question with a viable answer option?

Directrix (directrix):

A basketball goal has the equation x² + y² + 4x - 2y = 0. A ball with a radius of 1 comes down at the at the point of intersection of the lines with equations 7x - 4y = -15 and 2x + 3y = 4.

Directrix (directrix):

Which one of the following happens to the ball? a) Scores 2 points without touching the rim b) Scores 2 points but touches the rim c) Touches the rim but doesn't go through the hoop d) Misses hoop entirely e) Ball is too big to go through hoop

OpenStudy (campbell_st):

I think this question should have been written in a different manner that allows the student to step through the different parts of the question. (a) Show the centre of the circle is ... (b) Show the point of intersection of the lines.... (c) find the distance between the centre and point of intersection.. this way if the student makes a calculation error it doesn't result in a zero. by giving some of the information then a student may be able to complete some parts and be partially rewarded. Really the text of the question is asking to show fluency in several parts then asking you for an analysis based on the results. Personally I think it stinks.

OpenStudy (mathmate):

I do not think the question is pedagogically correct for a math question. 1. It expects students know the details of a basketball game, which is not a given for all students. It is like a probability question expecting a student to know what a full-house means. In my opinion, it should at least describe that a ball should pass through the hoop (goal ??) to score, and what the scoring system is, and under what conditions the ball is deemed to pass through (not touching the rim, or touching the rim, by how far) 2. It does not define or state that the x-y plane is horizontal, leaving a lot of interpretation on the student, who might score a zero even though he/she masters analytic geometry.